# Lost in Transcription, Found in Distribution Shift: Demystifying Hallucination in Speech Foundation Models

### v1

@@ Appendix A Appendix
A.1 Experiments
A.1.1 Datasets
@@
A.1.2 Perturbation
To evaluate the robustness of ASR models under varying conditions, we apply the following perturbations to the audio inputs:
  • White Noise: Gaussian noise is added at a low amplitude to simulate environmental interference.
  • Time Stretching: The audio is randomly stretched by a factor between 0.9 and 1.1, altering the speed while preserving pitch.
  • Pitch Shifting: The pitch is randomly shifted by up to ±2 semitones to mimic natural variations in speech.
  • None: No perturbation is applied, serving as the baseline for comparison.
These perturbations are designed to replicate real-world challenges such as background noise, speaker variability, and recording inconsistencies.
Type Description
None We do not apply any modification to the input speech. Serving as the baseline for comparison.
White Noise Gaussian noise is added at a low amplitude, simulating environmental interference.
Time Stretching The audio is randomly stretched by a factor between 0.9 and 1.1, altering the speed without affecting the pitch.
Pitch Shifting The pitch is randomly shifted by up to ±2 semitones.
Table 5: Perturbation details for synthetic shift.
A.1.3 Models
@@ A.2 Prompts
A.3 Results
We provide a detailed analysis of the experimental results, focusing on the highest and lowest performing models across the study. To offer a comprehensive overview, we present Hallucination Error Rate (HER) and Word Error Rate (WER) across domain shifts for all models, as shown in
@@ A.3 Results
. Additionally, we include fine-grained error analysis which highlights the differences between coarse-grained and fine-grained error categorization.
@@
We also calculate HER to WER ratio. As robust models would exhibit a smaller gap between hallucination and non-hallucination errors.
@@
Furthermore, we present the percentage of non-hallucination errors across datasets and models, categorizing them into Phonetic (P), Oscillation (O), and Language (L) errors. This analysis provides deeper insights into the types of errors that are most frequent and their distribution across different dataset-model combinations.
@@
We also highlight the overall distribution across all datasets and the robustness of both levels (coarsegrained and finegrained) in correctly identifying hallucination.
Key Findings:
  • The highest and lowest performing models exhibit significant variations in HER and WER under domain shifts, with some models showing robustness while others struggle.
  • Fine-grained error analysis reveals that certain error types (e.g., Oscillation) are more prevalent in specific dataset-model combinations.
  • Non-hallucination errors, particularly Phonetic and Language errors, dominate in certain scenarios, providing actionable insights for improving model performance.
These results underscore the importance of considering both hallucination and non-hallucination errors when evaluating ASR systems, as well as the need for domain-specific adaptations to enhance robustness.

### v2

@@ Appendix A Appendix
A.1 Hallucination
Table 9 illustrates the key differences between phonetic errors and hallucinations in ASR outputs. In the given example, the reference sentence "It’s hard to recognize speech" is misrecognized phonetically as "It’s hard to wreck a nice beach", demonstrating a typical pronunciation-based error where acoustically similar words are confused. In contrast, the hallucination example shows the model generating "Now move to me" instead of the reference "There are more coming", introducing semantically unrelated content not supported by the input speech. This highlights that phonetic errors stem from acoustic-perceptual confusions, while hallucinations involve the model producing arbitrary, contextually inconsistent text. The distinction is crucial for analyzing ASR failures, as each type requires different mitigation strategies—improving acoustic modeling for phonetic errors versus constraining language generation for hallucinations.
A.2 Experiments
In this section, we provide more details about our experimental setup.
A.2.1 Datasets
We provide a brief description of the datasets we use in our experiment in Table 10.
A.2.2 Models
We provide a brief description of the models we use in our experiment in Table 11.
A.3 Prompts
We show the prompts for coarse-grained and fine-grained error classification in Figures 5 and 6, respectively.
A.4 Results
We provide a detailed analysis of the experimental results, focusing on the highest and lowest performing models across the study. To offer a comprehensive overview, we present Hallucination Error Rate (HER) and Word Error Rate (WER) across domain shifts for all models, as shown in 7.
Additionally, we include fine-grained error analysis in Table 13, which highlights the differences between coarse-grained and fine-grained error categorization.
We also calculate the HER to WER ratio and present the numbers in Table 14. Robust models would exhibit a smaller gap between hallucination and non-hallucination errors.
We provide results for transducer-based models along with Canary-1B in Table 12.
Furthermore, we present the percentage of non-hallucination errors across datasets and models, categorizing them into Phonetic (P), Oscillation (O), and Language (L) errors in Table 15. This analysis provides deeper insights into the types of errors that are most frequent and their distribution across different dataset-model combinations.
We also highlight the overall distribution across all datasets and the robustness of both levels (coarse-grained and fine-grained) in correctly identifying hallucination in Figure 8. These results underscore the importance of considering both hallucination and non-hallucination errors when evaluating ASR systems, as well as the need for domain-specific adaptations to enhance robustness.
A.4.1 LLM Evaluation
We recognize that relying on GPT-4o for hallucination detection introduces the possibility that the evaluation model itself may produce incorrect labels, leading to false positives or negatives that could skew the assessment of ASR models. Given the length of our prompt (as shown in Figures 5 and 6), this also raises concerns about the “lost in the middle” effect Liu et al. (2023), which can further increase the risk of hallucination during evaluation.
To mitigate these challenges, our evaluation framework restricts GPT-4o’s output to only the final classification decision, explicitly avoiding chain-of-thought generation that could introduce errors Kim et al. (2024b). This is achieved through carefully designed prompts and a deterministic decoding strategy (temperature=0.0), ensuring consistent and reproducible outputs. Consequently, GPT-4o-mini consistently produces only the classification decision, minimizing the likelihood of introducing evaluation-induced hallucinations.
Our findings show strong alignment between GPT-4o-mini’s evaluations and human judgments, with minimal false positives or false negatives. While such errors are inevitable in any classifier short of a perfect oracle, they are substantially less frequent in our approach compared to heuristic-based methods, which often misclassify phonetic and oscillation errors as hallucinations. False negatives are particularly rare, and most false positives arise from phonetic or minor oscillation errors—very rarely (less than 1%) from semantically correct transcriptions.
Effect of Text Normalization
We also examine the impact of text normalization that we use to post-process the text for error rate calculation and hallucination detection. While normalization removes filler words and disfluencies—potentially important for applications requiring verbatim transcripts Ma et al. (2023); Talafha et al. (2023)—we find that it has negligible impact on hallucination classification. We report the results in Table 6 and 7 for coarse-grained and fine-grained evaluations, respectively.
In our evaluation of 500 examples, we observe no transitions from “No Error” to “Hallucination Error,” and only about a 1% transition from “Hallucination Error” to “Non-Hallucination Error” across both coarse-grained and fine-grained evaluations.
From (Normalized) To (Orthographic) Count
HE NE 2
HE NHE 7
NE NHE 1
NHE HE 2
Table 6: Transitions from normalized to orthographic transcriptions (coarse-grained) in our 500-example evaluation. Abbreviations: HE - Hallucination Error, NE - No Error, NHE - Non-Hallucination Error.
From (Normalized) To (Orthographic) Count
HE NE 1
HE OE 1
HE PE 5
HE HE 1
OE HE 1
PE HE 2
PE LE 1
Table 7: Transitions from normalized to orthographic transcriptions (fine-grained) in our 500-example evaluation. Abbreviations: HE - Hallucination Error, NE - No Error, OE - Oscillation Error, PE - Phonetic Error, LE - Language Error.
The overall agreement scores of 0.9324 (coarse-grained) and 0.9495 (fine-grained) further underscore the robustness of our evaluation framework, confirming that normalization does not compromise the integrity of hallucination detection. These findings demonstrate that for our use case, normalization preserves semantic fidelity without introducing artifacts that would affect hallucination classification, ensuring that our evaluation remains consistent and reliable even when “disfluencies“ are removed.
Cost Analysis
We acknowledge that using proprietary models such as GPT-4o-mini incurs significant costs, which may limit scalability. To address this concern, we conduct a detailed cost analysis comparing our LLM-based evaluation framework to a human evaluation baseline, which serves as an upper bound for cost. This analysis scales to 1M segments, each approximately 20 seconds in duration, totaling roughly 5,000 hours.
For the LLM-based evaluation, the combined fine- and coarse-grained prompts take approximately 1,000 tokens per example (upper bound), with most input tokens cached. The error class outputs typically require fewer than five tokens on average. At the time of evaluation, the pricing is \$0.075 per million input tokens and \$0.600 per million output tokens. Scaling this setup to 1M segments:
  • Input cost: 1, [^0] M tokens $\times$ \$0.075/1M = \$75.
  • Output cost: 5M tokens $\times$ \$0.600/1M = \$3.
The total upper-bound cost is therefore \$78. With batch inference, we expect this could be reduced by 50%, making it even more cost-effective.
In contrast, our human evaluation baseline requires approximately 10 minutes to evaluate 50 examples. Scaling to 1M examples:
  • Human evaluation time: (10 minutes / 50 examples) $\times$ 1M = 200,000 minutes ($\approx$ 3,333 hours).
  • Human evaluation cost (at \$10/hour): $\approx$ \$30,000.
Overall, this analysis demonstrates that our LLM-based evaluation framework is approximately 375 times cheaper than human evaluation, underscoring its scalability and cost-effectiveness for large-scale evaluation tasks.
A.4.2 Synthetic Shift
We provide results for the additional synthetic perturbations in Table 8.
Model Cross-lingual Noise Distortion Echo Pub Noise Reverberation
WER CER cHER fHER WER CER cHER fHER WER CER cHER fHER WER CER cHER fHER WER CER cHER fHER
distil-medium.en 25.95 18.06 14.60 14.30 20.89 14.26 8.50 7.80 32.42 20.74 20.90 19.00 41.49 28.10 30.00 28.20 23.98 16.00 11.00 10.60
whisper-large-v3 22.93 16.43 12.30 10.90 17.71 12.79 6.40 6.40 32.74 24.14 12.80 12.10 36.31 27.24 23.60 23.10 21.21 15.16 8.30 9.20
wav2vec2-large-xlsr-53-english 53.33 35.74 25.60 28.80 34.65 18.69 9.90 9.60 83.62 53.84 39.50 47.10 52.61 31.51 30.10 28.10 42.38 24.41 13.50 15.90
Table 8: WER, CER, cHER, and fHER of different models evaluated across various perturbation scenarios: Cross-lingual Noise, Distortion, Echo, Pub Noise, and Reverberation. Abbreviations: WER/CER - Word/Character Error Rate; cHER - Coarse-grained Hallucination Error Rate; fHER - Fine-grained Hallucination Error Rate.
Error Type Reference Hypothesis
Phonetic It’s hard to recognize speech It’s hard to wreck a nice beach
Hallucination There are more coming Now move to me
Language ripped ocean jumper Rips Ocean Jumper
Oscillation oh so it’s uh yeah Oh, I see, I see, I see, I see, I see, I see,
Table 9: Examples of different ASR error types, including phonetic confusions, hallucinations, language errors, and oscillation artifacts.
@@ A.4.2 Synthetic Shift
Model BERSt Primock57 Adversarial Salt-multispeaker-eng Supreme-court-speech
WER CER cHER fHER WER CER cHER fHER WER CER cHER fHER WER CER cHER fHER WER CER cHER fHER
canary-1b 37.4 37.4 18.8 14.7 18.0 17.2 6.0 6.6 27.9 27.5 46.3 41.6 3.3 3.3 2.1 1.6 16.0 15.7 18.9 21.3
parakeet-rnnt-1.1b 26.1 27.7 16.4 13.9 17.0 16.2 5.7 6.3 17.5 17.2 29.8 30.6 3.1 3.1 1.6 2.1 15.5 15.3 17.4 18.8
parakeet-tdt-1.1b 27.7 28.6 12.6 11.1 15.8 15.0 5.5 6.4 19.4 19.1 32.2 30.2 3.3 3.3 1.1 1.6 15.7 15.5 15.7 19.1
Table 12: WER, CER, cHER, and fHER for different models across datasets.